OpenStudy (anonymous):

Write each function in vertex form. y=a(x-h)^+k where h= -b/2a y=x^-6x+9 Please help me step by step please

OpenStudy (anonymous):

@Psymon please please help me

OpenStudy (psymon):

Do you know how to complete the square?

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

Can you show me, I have 4 more similar questions like this

OpenStudy (psymon):

Ah, okay. Yeah, this requires that you know how to complete thwe square.

OpenStudy (anonymous):

Okay, I think I know what that is, just maybe not the name of it.

OpenStudy (psymon):

Ill show you a quick example and see if it rings a bell: x^2 + 4x - 19 (x^2 + 4x + ___) = 19 + _____ (x^2 + 4x + 4) = 19 + 4 (x+2)^2 = 23

OpenStudy (anonymous):

the ^ is confusing me, can u just write the exponent please

OpenStudy (psymon):

Just write the exponent?

OpenStudy (anonymous):

I have no clue what you did up there basically, and how does that have to do with the question up there? :/

OpenStudy (psymon):

I wanted to see if the process looked familiar to you, because its whats required for this problem. It has nothing todo with the problem other than to see if youve seen it before.

OpenStudy (anonymous):

Then no i'm sorry it wasn't familiar

OpenStudy (psymon):

Alrighty. So, you know what a perfect square is?

OpenStudy (anonymous):

Yes a number multiplied by its self

OpenStudy (anonymous):

can you please show me how to do this one so I can do the other 4 please, its really late here

OpenStudy (psymon):

Right. Well, the same thing exists with polynomials. Exceot a perfect square polynomial would be a factor multiplied by itself. For example (x+1)^2. So when you see that form you want a(x-h)^2 +k  That portion is always a perfect square. Problem is when we're given a problem like yours, x^2-6x + 9, we need to see if its a perfect square or not. If it isnt a perfect square, we need to turn it into one. Thats what completing the square does, it find s way to force our problem to become a perfect square and look like vertex form